OpenStudy (anonymous):

When the quadratic equation y = (x − 5)(2x + 3) is written in standard form, what is the value of the coefficient "b"? answers: 7 −7 2 −15

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

\[2x ^{2}-7x-15\]

OpenStudy (anonymous):

foil

OpenStudy (anonymous):

ok. can you help me with another problem?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Identify which equation is a quadratic equation. y = −5x − 3 y = −4x(x + 10) + 34 y + x2 = (x − 4)(x + 3) y − 2 = (x − 2)(x2 + 1)

OpenStudy (anonymous):

quadratic means that a variableis raised to the power of 2

OpenStudy (anonymous):

ok so the second one?

OpenStudy (anonymous):

no you have to simplify them

OpenStudy (anonymous):

there may be more than one

OpenStudy (anonymous):

ok. thank you